If the roots of a cubic equation are: x=-5/2, x=1/4 & x=3. Determine this cubic expression by expansion, with these roots.
Answer by stanbon(75887)

If the roots of a cubic equation are: x=-5/2, x=1/4 & x=3. Determine this cubic expression by expansion, with these roots.
--------------
If those numbers are roots, then:
x+(5/2)=0
or
x-(1/4)=0
or
x-3=0
-------------
Which means:
2x+5=0
or
4x-1=0
or
x-3=0
------------
Taking the product of these you get:
(2x+5)(4x-1)(x-3)=0
(2x^2-2x+20x-5)(x-3)=0
2x^3-2x^2+20x^2-5x-6x^2+6x-60x+15=0
2x^3+12x^2-59x+15=0
----------
Cheers,
Stan H.